If 2x - 6 = 4, then what is the value of x?
Found 2 solutions by Alan3354, lenny460:
Answer by Alan3354(69443)

If 2x - 6 = 4, then what is the value of x?
--------------
Add 6
Divide by 2

Answer by lenny460(1073)

If 2x - 6 = 4, then what is the value of x?
2x - 6 = 4
Add 6 to each side of the equation.
2x - 6 + 6 = 4 + 6
Combine like terms:
-6 + 6 = 0
4 + 6 = 10
Therefore:
2x = 10
Divide each side by 2
2x/2 = 10/2
x = 5
The Answer:
x = 5



Lennox Obuong
Algebra Student
Email: obuong3@aol.com